OpenStudy (anonymous):

Task 3 Part 1: Create a scenario for an arithmetic sequence. For example: Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. __________________________________________________________ Part 2: Create a scenario for a geometric sequence. For example: Anthony goes to the gym for ______ minutes on Monday. Every day he _________his gym time by ____________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? __________________________________________________________ Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

I need someone who will explain this step by step

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

If you can't explain can you pls check my answers?

OpenStudy (anonymous):

Jasmine practices the piano for _10_____ minutes on Monday. Every day she ____increases_______ her practice time by __15 minutes_______. If she continues this pattern, how many minutes will she practice on the 7th day? Use your scenario to write the function for the 7th term in your sequence using sequence notation. Monday- 10 minutes Tuesday - 25 minutes Wednesday - 40 minutes Thursday - 55 minutes Friday - 70 minutes Saturday - 85 minutes Sunday - 100 minutes She will practice piano for 100 minutes on the 7th day, Sunday. Function: An = a1 + d(n – 1) Day N = 10 +15(n-1) Day 7= 10+15 (7-1) Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10%________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. (I rounded the decimals here) Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes He will spend 29.282 minutes at the gym on the 5th day, Friday. Sn=a*1−r^n/1−r where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. S(5)=20 1-1.1^5 / 1-1.1 =122.102

OpenStudy (anonymous):

Is this right? If not, pls explain why.

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (anonymous):

I'll give a medal if you help

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Do you think I'm right @Tommymair

OpenStudy (anonymous):

hard to read all of this, but the first one looks good

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think its good

OpenStudy (anonymous):

Starting from Anthony is part 2

OpenStudy (anonymous):

second one too maybe shouldn't round but maybe should

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I deleted what I had before rounding. Probably shouldn't have

OpenStudy (anonymous):

monday \(20\) tuesday \(20\times 1.1=22\) wednesday \(20\times 1.1^2=24.2\) tec

OpenStudy (anonymous):

I'm a little confused, my older sister kind of helped me with this one, but didn't give much of an explanation.

OpenStudy (anonymous):

So @satellite73 is this right?

OpenStudy (anonymous):

got any idea on part 3?

OpenStudy (anonymous):

I don't even know if part 1 & 2 r right

OpenStudy (anonymous):

eh whatever

OpenStudy (anonymous):

i have till jan 24 to finish both segments or ill get kicked out of school

OpenStudy (anonymous):

so im willing to take some risks

OpenStudy (anonymous):

Me 2, but I need an A, this is my 20th week!!!

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

oh i don't

OpenStudy (anonymous):

i can get whatever grade

OpenStudy (anonymous):

I'm trying to get into an Ivy League College

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

what school do you go to?

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

that sounded creepy

OpenStudy (anonymous):

Ha ha

OpenStudy (anonymous):

\[\Large S _{n} = a \frac{ 1 - r^n }{ 1-r }\]

OpenStudy (anonymous):

a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days.

OpenStudy (anonymous):

add up the times for the 5 days and see if they match.

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

sorry i had to do something else

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

what are we doing?

OpenStudy (anonymous):

Part 3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i can't help with that sorry

OpenStudy (anonymous):

im horrible at algebra

OpenStudy (anonymous):

Part 3

OpenStudy (anonymous):

pls at least try

OpenStudy (anonymous):

ill try but theres no guarantee

OpenStudy (anonymous):

what do you want me to do?

OpenStudy (anonymous):

\[s(5)= 20 \times \frac{ ( 1 - 1.1^5 ) }{ ( 1 - 1.1) } = ?\]

OpenStudy (anonymous):

Solve this, if you want use a calculator on Google, but include parenthesis

OpenStudy (anonymous):

I got 122.102

OpenStudy (anonymous):

U?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

I don't want to do all of the work

OpenStudy (anonymous):

I'm sorry i keep on closing the tab

OpenStudy (anonymous):

then i have to find the question again

OpenStudy (anonymous):

Whatever, I already did ALL of it

OpenStudy (anonymous):

Task 3 Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for _10_____ minutes on Monday. Every day she ____increases_______ her practice time by __15 minutes_______. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. Monday- 10 minutes Tuesday - 25 minutes Wednesday - 40 minutes Thursday - 55 minutes Friday - 70 minutes Saturday - 85 minutes Sunday - 100 minutes She will practice piano for 100 minutes on the 7th day, Sunday. function : An = a1 + d(n – 1) Day N = 10 +15(n-1) Day 7= 10+15 (7-1) part 2 Create a scenario for a geometric sequence. For example, Anthony goes to the gym for __20____ minutes on Monday. Every day he ___increases______his gym time by ____10%________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Be sure to fill in the blanks with the words that will create a geometric sequence. Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. (i rounded the decimals here) Mon - 20 minutes Tue - 22 minutes Wed - 24.2 minutes Thu - 26.62 minutes Fri - 29.282 minutes He will spend 29.282 minutes at the gym on the 5th day, Friday. part 3 Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario. Sn=a*1−r^n/1−r where a is the first term and r is the common ratio. Here a = 20 minutes and r = 1.1 and to find out the total time at the gym from Mon-Fri, n = 5 days. S(5)=20 1-1.1^5 / 1-1.1 =122.102

OpenStudy (anonymous):

holly

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

Ur welcome

OpenStudy (anonymous):

Pls fan me, I'll fan you. Maybe we can work on other stuff together some time since we're in the same class

OpenStudy (anonymous):

aight

OpenStudy (anonymous):

what grade do you have in the class?

OpenStudy (anonymous):

Wait, let me check

OpenStudy (anonymous):

I have 90.85 (rounded to 91)

OpenStudy (anonymous):

But I'm in 7th grade taking high school course, so don't judge me, pls

OpenStudy (anonymous):

U?

OpenStudy (anonymous):

i have a 95

OpenStudy (anonymous):

but thats good tho

OpenStudy (anonymous):

Awesome!!! Congrats!!!

OpenStudy (anonymous):

im in 7th grade too

OpenStudy (anonymous):

Thx!!!

OpenStudy (anonymous):

and i go to a high school

OpenStudy (anonymous):

well technically i don't

OpenStudy (anonymous):

I don't go to high school, but I do online high school math. That's the only high school course I take

OpenStudy (anonymous):

i go to a highshool that has a 7th and 8th grade

OpenStudy (anonymous):

But I also do this thing called Kumon, it's annoying

OpenStudy (anonymous):

Oh,

OpenStudy (anonymous):

whats kumon?

OpenStudy (anonymous):

I'm sorry, but I have a sixth sense of detecting liars. Are u lying to me. Don't be offended if I'm wrong! I just have the same feeling I usually have when someone is lying.

OpenStudy (anonymous):

Kumon is for gifted kids

OpenStudy (anonymous):

lying of what?

OpenStudy (anonymous):

i go to MAST Academy

OpenStudy (anonymous):

Idk, in general, anything

OpenStudy (anonymous):

6th best highschool in FL

OpenStudy (anonymous):

well i have a 94.86 not a 95

OpenStudy (anonymous):

It's just that you said ur grade didn't matter and if it doesn't how do you have a 95

OpenStudy (anonymous):

as long as i complete it I'm fine

OpenStudy (anonymous):

but if i can get a high grade then better

OpenStudy (anonymous):

U made it seem like u didn't really care much about school and u were a slacker, that's also why I wondered why ur taking an honors class because this project is for honor students

OpenStudy (anonymous):

i have to go

OpenStudy (anonymous):

But, whatever, I was wrong.